OpenStudy (anonymous):

If the circle (x + 5)2 + (y - 4)2 = 35 is translated 5 units down and 4 units right, then the sum of the new y-intercepts is?

OpenStudy (anonymous):

If the circle (x + 5)^2 + (y - 4)^2 = 35 is translated 5 units down and 4 units right, then the sum of the new y-intercepts is

OpenStudy (binary3i):

(x +5)^2 + (y-4)^2= 35 new position, it becomes. (x+5-4)^2 +(y-4+5)^2=35 (x+1)^2 + (y+1)^2 =35 now for y intersept put x=0 1+ (y+1)^2 = 35 (Y+1)^2=34 \[(y+1)= \pm \sqrt{34}\] \[y=\pm \sqrt{34} -1\] sum of intersepts. \[- \sqrt{34} +1 + \sqrt{34} +1=2\]

OpenStudy (binary3i):

sorry -2

OpenStudy (anonymous):

thank you

OpenStudy (binary3i):

welcom